OpenStudy (anonymous):

The radius of a cylindrical popcorn tin is 3x+1 in. The height of the tin in 3 times the radius. What is the surface area of the cylinder? Write your answer in simplified form.

OpenStudy (marigirl):

have u gt the final ans by any chance?

OpenStudy (radar):

Does the surface include the lid......or is there a lid?

OpenStudy (marigirl):

??

OpenStudy (radar):

R= 3x + 1 then Diameter D = 6x + 2 or 2R Circumference C= pi(6x + 2) Height H=3r or 9x + 3. The cylinder (minus bottom & lid) area A=Ch=pi(6x + 2)(9x + 3) =pi(54x^2 + 36x + 6) Now the area of the bottom is pir^2 or pi(3x + 1)^2 and so would be the lid if there was one. Just add them.

OpenStudy (radar):